# Adversarial Training: Enhancing Out-of-Distribution Generalization for Learning Wireless Resource Allocation

### v1

@@ IV Case Study: Optimizing Hybrid Precoding
In this task, ${\bf H}=[{\bf h}_{1},\cdots{\bf h}_{K}]^{T}\in\mathbb{C}^{K\times N_{T}}$ is the environment parameter. To adapt to a wide range of transmit power and channel strength, $S\!N\!R\triangleq\left\|{\bf H}\right\|_{F}^{2}P_{tot}/\sigma^{2}$ is also regarded as an environment parameter. Then, the DNN for learning the hybrid precoding is $({\bf W}_{RF},{\bf W}_{BB})=f_{\bf y}({\bf H}/\!\left\|{\bf H}\right\|_{F},S\!N\!R;{\bm{\theta}})$ with two inputs, where the channel is normalized to avoid redundancy in the input information. Meanwhile, since both (15b) and (15c) are unrelated to ${\bf H}$ and can be guaranteed by the projection method in [15], only the $K$ QoS constraints in (15d) need to be satisfied using a multiplier network denoted as $f_{\bm{\lambda}}({\bf H}/\!\left\|{\bf H}\right\|_{F},S\!N\!R;{\bm{\xi}})$.
When applying UAT, we only need to generate adversarial examples for ${\bf H}/\!\left\|{\bf H}\right\|_{F}$, since $S\!N\!R$ is a scalar with finite range that does not need to be generalized. To ensure that the adversarial examples are normalized, we modify (14) as
@@ V Simulation Results
We use the following two metrics to evaluate performance. Average SR (ASR) is the averaged SR over a test set, where the SR of a sample is recorded as zero if the constraints are not satisfied. Violation Ratio (VR) is the proportion of samples in a test set that violate the constraints.
To obtain a benchmark for comparison, we simulate a difference-of-convex (DC) programming-based numerical algorithm [14] to optimize hybrid precoders. Instead of selecting from the array responses computed by the angles of departure (AoDs) of all propagation paths [14], we set ${\bf W}_{RF}$ as the phases of the first $N_{RF}$ right singular vectors of ${\bf H}$.
In simulations, the BS equipped with $N_{T}=16$ antennas and $N_{RF}=6$ RF chains serves $K=4$ users with data-rate requirements of $\gamma_{1}=\cdots=\gamma_{K}=1$ bps/Hz. $P_{tot}/\sigma^{2}$ is uniformly selected from the range of 0$\sim$20 dB during training, and is set as 10 dB for testing unless otherwise specified.
We compare UAT with two existing AT methods in [11, 10]. They are also incorporated into PDL but generate adversarial examples in other ways. Following the configurations in [11, 10], the number of descent steps is 7, step size is 2, and projection radius is 8. For UAT, the tuned hyper-parameters are $E_{A}=50$, $T=5$, $\epsilon=5$. The PDL in [2] without any AT serves as a baseline.
Both $f_{\bf y}(\cdot\,;{\bm{\theta}})\!$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})\!$ are realized by three kinds of DNNs designed for hybrid precoding: convolutional neural networks (CNN) [16], three-dimensional graph neural network (3D-GNN) [15], and recursive GNN (RGNN) [17]. The optimizer is Adam, $\eta_{1}=\eta_{2}=0.001$, $N_{B}=100$, batch normalization is used for CNN and 3D-GNN, activation function is ${\rm Tanh}(\cdot)$ for RGNN and ${\rm ReLU}(\cdot)$ for CNN and 3D-GNN. Other hyper-parameters are listed in Table I, where $3\times 3$ convolution kernels are used in the hidden layers of CNN followed by a fully-connected layer with 512 neurons.
TABLE I: Fine-tuned Hyper-Parameters
@@
These DNNs are trained with 100,000 samples of Rayleigh fading channel (i.e., ${\bf h}_{k}\sim\mathcal{CN}({\bf 0},{\bf I}_{N_{T}})$). Once trained, they are not fine-tuned, in order for examining their OOD generalization performance.
@@ V-A Rician Fading Channel
The OOD test samples are generated from the following Rician fading channel model
${\bf h}_{k}=\sqrt{{\kappa}/({\kappa+1})}{\bf a}(\phi_{k})+\sqrt{{1}/({\kappa+1})}{\bf z}_{k},$ (17)
where ${\bf a}(\phi_{k})=[1,{\rm e}^{{\rm i}\pi\sin(\phi_{k})},\cdots,{\rm e}^{{\rm i}(N_{T}-1)\pi\sin(\phi_{k})}]^{T}$ is the array response of the line-of-sight (LOS) path with AoD $\phi_{k}\sim\mathcal{U}(0,2\pi)$, ${\bf z}_{k}\sim\mathcal{CN}({\bf 0},{\bf I}_{N_{T}})$, and $\kappa$ is the Rician factor.
In Fig. 1, we provide the performance of training RGNN by PDL or UAT, with legend PDL(ID) or UAT(ID) when tested in Rayleigh channel and legend PDL(OOD) or UAT(OOD) when tested in Rician channel. Fig. LABEL:sub@fig:epoch-a shows that RGNN trained by both methods can achieve comparable ASR to DC in Rayleigh channel. When tested in Rician channel, the RGNN trained by UAT attains a 4.7% higher ASR (normalized by DC) than the RGNN trained by PDL. Fig. LABEL:sub@fig:epoch-b shows that DC can satisfy QoS constraints for all samples on both channels. When tested in Rayleigh channel, RGNN trained by both methods exhibits extremely low VR (about 0.3%). When tested in Rician channels, the RGNN trained by UAT has a 4.9% lower VR than that trained by PDL, thanks to the consistent objective in (15a) and constraints in (15d), even though no constraint satisfaction is considered in the cost of our DRO formulation. Besides, the inference time of RGNN is 15.48 ms, but DC requires 5498 ms on the same CPU.
@@
In Fig. 2, we compare the OOD generalization performance of RGNN and 3D-GNN trained by different methods. When the distribution of test samples differs noticeably from Rayleigh channel (say $\kappa\geq 10$ dB), UAT improves OOD generalization by approximately 5% in both ASR and VR for the two GNNs. The ASR of UAT-trained DNNs still retains 85% of the ASR of DC even when $\kappa=20$ dB. The DNNs trained by the AT methods in [11, 10] exhibit marginal OOD generalization gains over PDL, due to not designed for unsupervised learning. Their performance degradation observed at $\kappa\leq 5$ dB (close to Rayleigh channels, i.e., ID samples) results from their overly broad perturbation range.
@@ V-B Spatially Correlated Channel
The OOD test samples are generated from the following spatially correlated channel model [18]
${\bf H}={\bf R}^{\frac{1}{2}}_{U}{\bf Z}{\bf R}^{\frac{1}{2}}_{A},$ (18)
where each element in ${\bf Z}$ is $z_{k,n}\sim\mathcal{CN}(0,1)$, ${\bf R}_{A}\in\mathbb{R}^{N_{T}\times N_{T}}$ is the antenna correlation matrix with element in the $i$th row and $j$th column being $\rho_{A}^{|i-j|}$, ${\bf R}_{U}\in\mathbb{R}^{K\times K}$ captures the inter-user correlation with diagonal elements being 1 and off-diagonal elements being $\rho_{U}$, and $\rho_{A},\rho_{U}\in[0,1]$.
In Table II, we provide the OOD generalization performance of the DNNs trained by PDL or UAT when they are tested on correlated channels. As shown by the values in boldface, the OOD generalization gains of UAT over PDL are more pronounced in highly correlated channels ($\rho_{U},\rho_{A}\geq 0.6$), which often arise in scenarios with densely distributed users, limited scatters, or closely spaced antennas. In particular, for RGNN, the OOD gain ranges from 4.1% (i.e., 99.7%--95.6%, shown in red) to 20.3% (shown in green, similarly hereafter), whereas for 3D-GNN, the gain ranges from 3.1% to 13.8%. The lower gain of 3D-GNN is attributed to its inherent relatively better OOD generalizability. For CNN, UAT significantly boosts ASR by 10.2$\sim$22.6across all values of $\rho_{U}$ and $\rho_{A}$, not limited to highly correlated channels. This is because UAT can mitigate the overfitting of CNN.
@@ V-C Sparse Channel and 3GPP Channel
Finally, the OOD test samples are generated from two millimeter-wave channel models, which are commonly used for hybrid precoding. The first is an $L$-path sparse channel,
${\bf h}_{k}=\sqrt{N_{T}/L}\textstyle\sum_{l=1}^{L}\alpha_{k,l}{\bf a}(\phi_{k,l}),$ (19)
where $\alpha_{k,l}\sim\mathcal{CN}(0,1)$ is a complex gain, and ${\bf a}(\phi_{k,l})$ is the array response for AoD $\phi_{k,l}\sim\mathcal{U}(0,2\pi)$.
@@
In Table III, we provide the OOD generalization performance of the DNNs trained by PDL or UAT when they are tested on the two kinds of channels. It is evident that UAT enhances the OOD generalizability of all three DNNs, even for the LOS and NLOS scenarios in 3GPP channel model. The results in boldface indicate that 3D-GNN outperforms both RGNN and CNN in terms of OOD generalization.
@@ VI Conclusions
In this letter, we reformulated AT for unsupervisedly training OOD-generalizable DNNs to optimize wireless resource allocation. We proposed the UAT method to solve the formulated AT problem, and applied the method to train three different DNNs for learning QoS-aware hybrid precoding. Simulation results showed that these DNNs trained by the proposed method can achieve superior OOD generalization on Rician fading, spatially correlated, sparse, and 3GPP channels, by being trained only on Rayleigh fading channels.

### v2

@@ IV Case Study: Optimizing Hybrid Precoding
[…]
When applying UAT, we only need to generate adversarial examples for ${\bf x}={\bf H}/\!\left\|{\bf H}\right\|_{F}$, since SNRS\!N\!R is a scalar with finite range that does not need to be generalized. To ensure that the adversarial examples are normalized, we modify (14) as
@@ V Simulation Results
We use the following two metrics to measure performance. Average SR (ASR) is the averaged SR over a test set, where the SR of a sample is recorded as zero if the constraints are not satisfied. Violation Ratio (VR) is the proportion of samples in a test set that violate the constraints.
To obtain a benchmark for comparison, we simulate a difference-of-convex (DC) programming-based numerical algorithm [20] to optimize hybrid precoders. Yet different from [20], we set ${\bf W}_{RF}$ as the phases of the first $N_{RF}$ right singular vectors of $\bf H$}, instead of selecting from the array responses computed by the angles of departure (AoDs) of all propagation paths that is only applicable to sparse channels.
[…]
We compare UAT with two existing AT methods in [18, 16] that generate adversarial examples for supervised learning. Following the configurations in [18, 16], the number of descent steps is 7, step size is 2, and projection radius is 8. For UAT, the fine-tuned hyper-parameters are $E_{A}=50$, $T=5$, $\epsilon=0.5$. The PDL in [1] without any AT serves as a baseline.
Both $f_{\bf y}(\cdot\,;{\bm{\theta}})\!$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})\!$ are realized by three kinds of DNNs designed for hybrid precoding: convolutional neural networks (CNN) [22], three-dimensional graph neural network (3D-GNN) [21], and recursive GNN (RGNN) [23]. The fine-tuned hyper-parameters are as follows. The optimizer is Adam, $\eta_{1}=\eta_{2}=0.001$, $N_{B}=100$, batch normalization is used for CNN and 3D-GNN, activation function is ${\rm Tanh}(\cdot)$ for RGNN and ${\rm ReLU}(\cdot)$ for CNN and 3D-GNN. Other fine-tuned hyper-parameters are listed in Table I, where $3\times 3$ convolution kernels are used in the hidden layers of CNN followed by a fully-connected layer with 512 neurons.
TABLE I: Tuned Hyper-Parameters for DNNs
@@
Since the UAT does not depend on data diversity, these DNNs are trained with 100,000 samples generated from a single distribution, where we take Rayleigh fading channel as an example. In fact, they can also be trained on other channel distributions and yield similar results. Once trained, these DNNs are not fine-tuned, in order to examine their OOD generalization performance.
@@ V-A Rician Fading Channel
The OOD test samples are generated from the Rician fading channel model: ${\bf h}_{k}=\sqrt{{\kappa}/({\kappa+1})}{\bf a}(\phi_{k})+\sqrt{{1}/({\kappa+1})}{\bf z}_{k}$, where ${\bf a}(\phi_{k})=[1,{\rm e}^{{\rm i}\pi\sin(\phi_{k})},\cdots,{\rm e}^{{\rm i}(N_{T}-1)\pi\sin(\phi_{k})}]^{T}$ is the array response of the line-of-sight (LOS) path with AoD $\phi_{k}\sim\mathcal{U}(0,2\pi)$, ${\bf z}_{k}\sim\mathcal{CN}({\bf 0},{\bf I}_{N_{T}})$, and $\kappa$ is the Rician factor.
In Fig. 1, we provide the performance of the RGNN trained by PDL or UAT (the results for the other two DNNs are similar and hence are not shown). It has legend PDL(ID) or UAT(ID) when tested in Rayleigh channel and it has legend PDL(OOD) or UAT(OOD) when tested in Rician channel. The Rician channel with $\kappa=10$ dB is a representative small-scale fading of LOS channels in UMa and UMi scenarios [24]. Fig. 1 (a) shows that RGNN trained by both methods can achieve comparable ASR to DC in Rayleigh channel. When tested in Rician channel, the RGNN trained by UAT attains a 4.7% higher ASR (normalized by DC) than the RGNN trained by PDL. Fig. 1 (b) shows that DC can satisfy QoS constraints for all samples on both channels. When tested in Rayleigh channel, RGNN trained by both methods exhibits extremely low VR (about 0.3%). When tested in Rician channels, the RGNN trained by UAT has a 4.9% lower VR than that trained by PDL, because the objective in (15a) and the constraints in (15d) are positively correlated, even though no constraint satisfaction is considered in the cost of our DRO formulation.
Besides, the inference time of RGNN is 15.48 ms, but DC requires 5498 ms on the same CPU, validating the advantage of learning-based approach for low decision latency.
@@
In Fig. 2, we compare the OOD generalization performance of RGNN and 3D-GNN trained by different methods. We can see that UAT improves OOD generalization by approximately 5% in both ASR and VR for the two GNNs when the distribution of test samples differs noticeably from Rayleigh channel (say $\kappa\geq 10$ dB). The ASR of UAT-trained DNNs still retains 85% of the ASR of DC even when $\kappa=20$ dB. The DNNs trained by the AT methods in [18, 16] exhibit marginal OOD generalization gains over PDL, due to not designed for unsupervised learning. Their performance degradation observed at $\kappa\leq 5$ dB (close to Rayleigh channels, i.e., ID samples) results from their overly broad perturbation range.
@@ V-B Spatially Correlated Channel
The OOD test samples are generated from the following spatially correlated channel model [25]: ${\bf H}={\bf R}^{\frac{1}{2}}_{U}{\bf Z}{\bf R}^{\frac{1}{2}}_{A}$, where each element in $\bf Z$} is $z_{k,n}\sim\mathcal{CN}(0,1)$, ${\bf R}_{A}\in\mathbb{R}^{N_{T}\times N_{T}}$ is the antenna correlation matrix with element in the $i$th row and $j$th column being $\rho_{A}^{|i-j|}$, ${\bf R}_{U}\in\mathbb{R}^{K\times K}$ captures the inter-user correlation with diagonal elements being 1 and off-diagonal elements being $\rho_{U}$, and $\rho_{A},\rho_{U}\in[0,1]$.
@@ V-C Sparse Channel and 3GPP Channel
Finally, the OOD test samples are generated from two millimeter-wave channel models, which are commonly used for hybrid precoding.
@@
In Table III, we provide the OOD generalization performance of the DNNs trained by PDL or UAT when they are tested on the two kinds of channels. It is evident that UAT enhances the OOD generalizability of all three DNNs, under all considered channels for testing. The results in boldface indicate that 3D-GNN outperforms both RGNN and CNN in terms of OOD generalization.
@@ VI Conclusions
In this paper, we reformulated AT problem for training OOD-generalizable DNNs in the unsupervised manner to optimize wireless resource allocation, and proposed an algorithm to solve the problem. We train three DNNs for learning QoS-aware hybrid precoding with the proposed UAT as a case study for a proof-of-concept demonstration. Simulation results showed that these DNNs trained by the proposed UAT can achieve superior OOD performance on Rician fading, spatially correlated, sparse, and 3GPP channels, despite being trained only on Rayleigh fading channels. In future works, we will evaluate the OOD generalization performance of the UAT over more realistic channels, such as real-world datasets.